# Kolmogorov-Arnold Attention: Is Learnable Attention Better For Vision Transformers?

### v2

@@ Abstract
Kolmogorov-Arnold networks (KANs) are a remarkable innovation consisting of learnable activation functions with the potential to capture more complex relationships from data. Presently, KANs are deployed by replacing multilayer perceptrons (MLPs) in deep networks, including advanced architectures such as vision Transformers (ViTs). Given the success of replacing MLP with KAN, this work designs and asks whether a similar replacement in the attention can bring benefits. In this paper, we design the first learnable attention called Kolmogorov-Arnold Attention (KArAt) for ViTs that can operate on any basis, ranging from Fourier, Wavelets, Splines, to Rational Functions. However, learnable activations in the attention cause a memory explosion. To remedy this, we propose a modular version of KArAt that uses a low-rank approximation. By adopting the Fourier basis into this, Fourier-KArAt and its variants, in some cases, outperform their traditional softmax counterparts, or show comparable performance on CIFAR-10, CIFAR-100, and ImageNet-1K datasets. We also deploy Fourier KArAt to ConViT and Swin-Transformer, and use it in detection and segmentation with ViT-Det. We dissect these architectures’ performance on the classification task by analyzing their loss landscapes, weight distributions, optimizer path, attention visualization, and transferability to other datasets, and contrast them with vanilla ViTs. KArAt’s learnable activation shows a better attention score across all ViTs, indicating better token-to-token interactions, contributing to better inference. Still, its generalizability does not scale with larger ViTs. However, many factors, including the present computing interface, affect the relative performance of parameter- and memory-heavy KArAts. We note that the goal of this paper is not to produce efficient attention or challenge the traditional activations; by designing KArAt, we are the first to show that attention can be learned and encourage researchers to explore KArAt in conjunction with more advanced architectures that require a careful understanding of learnable activations. Our open-source code and implementation details are available on: subhajitmaity.me/KArAt
@@ 1 Introduction
Figure 1: (a) Model parameters vs. Top-1 Test accuracy in ImegeNet-1K training of vanilla ViTs [24], Vision KAN (DeiT+KAN) by [17], ViT+KAN and Kolmogorov-Arnold Transformer (KAT) by [80].(b-i) The traditional softmax attention. (b-ii) The Kolmogorov-Arnold Attention (KArAt) replaces the softmax with a learnable operator, $\Phi^{i,j}$. (b-iii) Regular KArAt uses an operator matrix, $\Phi^{i,j}$ with $N^{2}$ learnable units acting on each row of ${\cal A}^{i,j}$, and is prohibitively expensive. (b-iv) Modular KArAt uses an operator $\widehat{\Phi}^{i,j}\in\mathbb{R}^{N\times r}$ with $r\ll N$, followed by a learnable linear projector $W\in\mathbb{R}^{r\times N}$.
@@
Artificial general intelligence has become a rapidly growing research direction, and Kolmogorov-Arnold Network (KAN) [50] marks a remarkable innovation in that. KANs with learnable activation functions can potentially capture more complex relationships and facilitate meaningful interaction between the model and human intuition. After training a KAN on a specific problem, researchers can extract the learned univariate functions that the model uses to approximate complex multivariable functions. By studying these learned functions, researchers can gain insights into the underlying relationships from the data and refine the model. KANs exhibit state-of-the-art performance in finding symbolic function representations [81], continual learning of one-dimensional functions [51].
@@
KANs were integrated with neural network architectures, the primary of them being conventional MLPs or convolution neural networks (CNNs). E.g., [28, 9, 25, 1, 75] combine KAN with CNNs, [44] combine KAN with U-Net, etc. Interestingly, [81] claimed to make the first fairer comparison between KANs and MLPs on multiple ML tasks on small-scale datasets. By setting the same parameter counts in both networks, [81] concluded that KANs with B-Spline basis significantly underperform compared to their MLP counterparts in all ML tasks except symbolic formula representation; [5] reinstates a similar claim for vision. However, KAN’s exploration of more advanced architectures, such as Transformers, remains limited. [17] replace the MLP layers inside the encoder blocks of data-efficient image Transformers (DeiTs) [73] with KANs and proposed DeiT+KAN, [80] proposed two variants: ViT+KAN that replaces the MLP layers inside ViT’s encoder blocks, and Kolmogorov-Arnold Transformer (KAT), albeit similar to ViT+KAN but with a refined group-KAN strategy.
@@
While KANs were claimed to be not superior to simple MLPs in vision tasks [81, 5], we were curious to find out how ViTs can see images through the lens of KANs. In that effort, Figure 1(a) demonstrates the performance of vanilla ViTs (-Tiny, -Small [76, 68], and -Base [24]) and their KAN counterparts on ImageNet-1K [21] training. While different variants of DeiT+KAN [17] and ViT+KAN [80] show a 5%–18% drop in the Top-1 accuracy compared to vanilla ViTs and DeiTs in ImageNet-1K, the KATs [80] show approximately 1%–2.5% gain in Top-1 accuracy compared to vanilla ViTs and DeiTs by keeping about the same number of parameters. These results show KANs require further investigation in more advanced architectures such as Transformers. Nevertheless, first, we want to understand why there is a discrepancy in these two seemingly equivalent implementations.
@@
In search of an answer, we realized that while DeiT+KAN and ViT+KAN replace the MLP layer with KAN in the Transformer’s encoder block, KAT implements a sophisticated group-KAN strategy that reuses a learnable function among a group of units in the same layer and chooses different bases for different encoder blocks. This strategy helps improve the accuracy by keeping almost similar parameter counts. Therefore, simply replacing MLPs with KANs might not guarantee better performance, but a properly designed KAN could. This perspective on the KANs’ performance in conjunction with Transformers opens many research questions. Importantly, we also note that the community remains put from designing a learnable multihead attention module, the heart of the Transformers. Therefore, in the rise of second-generation Transformers, such as Google’s TITAN [6] and SAKANA AI’s Transformer2 [70] that mimic the human brain, we want to ask: Is it worth deploying learnable multi-head self-attention to the (vision) Transformers?
@@
Finally, we note that there is a line of research that proposes efficient attention mechanisms in terms of sparse attention [63, 82, 67, 39, 85, 83, 31] or linear/kernelized attention [32, 58, 57, 56, 52, 55], to remedy computational and memory complexities of attention calculation. In contrast, this paper tries to understand how learnable multi-head self-attention (MHSA) modules can perform over regular self-attention used in ViTs, a defining technology in vision in the last six years. Our contributions are:
@@
Designing a Learnable Attention Module for ViTs. Since language is significantly more complicated, less structured, and has a higher information entropy than an image, we limit this pilot study to the ViTs [24] for image classification. In §3, we propose a general learnable Kolmogorov-Arnold multi-head self-attention, or KArAt, for ViTs that can operate on any basis. Due to the memory explosion, any learnable MHSA operator matrix in ideal KArAt causes a training bottleneck. We propose a low-rank approximation of the operator matrix to reduce memory requirements to a feasible level, but leave the possibility of exploring alternative techniques. By adopting the Fourier basis into this modular version, we design Fourier KArAt and its variants; see §4 for details.
@@
Benchmarking, Evaluation, and Analysis (§56). We benchmark Fourier KArAt on CIFAR-10, CIFAR-100, [40] and ImageNet-1K [21] datasets and evaluate their performance against vanilla ViTs (ViT-Tiny, -Small [76, 68], and -Base [24]), and other popular ViTs (ConViT [20] and Swin-Transformer [49]). We use small-scale datasets, SVHN [54], Oxford Flowers 102 [59], and STL-10 [18] for understanding the transfer learning capability of the models. Additionally, we embed KArAt with ViT-Det [46] for the object detection and segmentation task on MS COCO [48]. We dissect KArAts’ performance and generalization capacity by analyzing their loss landscapes, weight distributions, optimizer path, and attention visualization, and compare them with traditional softmax attention in vanilla ViTs.
@@ 2 Background
In MLP, the nonlinear activation functions, $\Phi(\cdot)$, are fixed [7]. For supervised tasks, given a training dataset $D$ with $N$ elements of the form (input, ground-truth) pairs, $\{(x_{i},y^{\star}_{i})\}_{i=1}^{N}$, the loss, $\mathcal{L}(X,Y^{\star}|\mathcal{W})=d_{Y}(a^{[L]},Y^{\star})$ (metric induced by the space $\mathbb{R}^{d_{L}}$) is calculated in the forward pass. The layerwise weights, $\{(W_{l},b_{l})\}_{l\in[L]}$ are learned by minimizing $\mathcal{L}(X,Y^{\star}|\mathcal{W})$.
Kolmogorov-Arnold Network (KAN) [51] is a neural network involving learnable activations parametrized by a chosen set of basis functions defined on a set of grid points or knots. The idea for this network stems from the Kolmogorov-Arnold Representation Theorem.
@@
[38] For any multivariate continuous function, $f:[0,1]^{n}\to\mathbb{R}$, there exists a finite composition of continuous single-variable functions, $\phi_{q,p}:[0,1]\to\mathbb{R},\Phi_{q}:\mathbb{R}\to\mathbb{R}$ such that
$f(x)=f(x_{1},x_{2},\cdots,x_{n})=\sum_{q=1}^{2n+1}\Phi_{q}\left(\sum_{p=1}^{n}\phi_{q,p}(x_{p})\right).$
Theorem 1 describes an exact finite expression using 2 layers, but [51] generalized this representation to multi-layers via KAN; it is analogous to the Universal Approximation Theorem [34]. For input $x^{[0]}$, an $L$-layer KAN is given as $\text{KAN}(x^{[0]})=\Phi^{L}\circ\cdots\circ\Phi^{l}\circ\cdots\circ\Phi^{1}(x^{[0]})$. In each KAN layer, the activation function $\Phi$ is learnable, and ${\Phi^{[l]}=[\Phi_{ij}^{[l]}]}$ operates on ${x^{[l-1]}\in\mathbb{R}^{d_{l-1}}}$ to produce ${x^{[l]}\in\mathbb{R}^{d_{l}}},$ such that:
@@
Originally, [51] chose the B-spline basis functions. That is, the activation functions were defined to be $\phi(x)=w(b(x)+\text{spline}(x))$, where $b(x)=\texttt{SiLU}(x)=\frac{x}{1+e^{-x}}$ and $\text{spline}(x)=\sum\limits_{i}c_{i}B_{i}(x)$, with $B_{i}(\cdot)$ being one of the $k$-th degree B-spline basis functions parametrized by the $G$ grid points on a uniform grid, $[-I,I]$. In this case, the representation for each function is given as:
$\Phi_{ij}^{[l]}(x^{[l-1]}_{j})=w^{[l]}_{ij}(\texttt{SiLU}(x^{[l-1]}_{j})+\sum\limits_{m=1}^{G+k-1}c^{[l]}_{ijm}B_{ijm}(x^{[l-1]}_{j})).$ (1)
The weights ${\{([c^{[l]}_{ijm}],[w^{[l]}_{ij}])\}_{l=1}^{L}}$ are learned by minimizing the loss for a supervised learning task. The bases for KANs’ activations could be Fourier [53, 78], wavelet [10], fractals [80], etc.
Multi-head self-attention (MHSA) in ViTs [24]. The encoder-only vanilla ViT architecture is inspired by the Transformer proposed in [74]. For simplicity, we do not mention the details of the layer normalization and other technicalities. Our central focus is the MHSA architecture.
@@ 3 How Can We Design Learnable Attention?
Last year, we witnessed a surge in embedding KANs in different DNN architectures [29, 28, 44]. To our knowledge, the work closely related to ours is KAT, where [80] replaces the MLP layers in the ViTs with KAN layers; also, see [17]. We note that, for attentive graph neural networks (GNNs), [27] unifies the scoring functions of attentive GNNs and names it as Kolmogorov-Arnold Attention (KAA). However, this attention is orthogonal to our MHSA design; deploying learnable MHSA is complicated. In this Section, we design learnable multi-head self-attention (MHSA) module for ViTs that operate inside its encoder blocks.
@@
Projection onto the probability simplex. The softmax function acts on each row of ${\cal A}$ to create a probability distribution; the learnable attention does not guarantee that. To ensure that each row vector of $\Phi(\mathcal{A})$ lies on a probability simplex, we project them onto the $\ell_{1}$-unit ball. That is, for each attention matrix, ${\cal A}^{i,j}\in\mathbb{R}^{N\times N}$, and for each $k\in[N],$ we want to have ${\|\tilde{\sigma}\left[({\cal A}^{i,j}_{k,:})\right]\|_{1}=1}$ and $\tilde{\sigma}\left[({\cal A}^{i,j}_{k,l})\right]\geq 0$ for $l\in[N]$. We cast a sparse approximation problem, whose variants have been well-studied in the past decade and arise frequently in signal processing [12, 23, 26] and matrix approximation [8, 69]; see §4. Algorithm 1 provides an $\ell_{1}$-projection pseudocode; for each $k\in[N]$, setting ${y=\tilde{\sigma}\left[({\cal A}^{i,j}_{k,:})\right]}$ and $m=N$ in Algorithm 1, we obtain the projected attention vector.

### v3

@@ Abstract
Kolmogorov-Arnold networks (KANs) are a remarkable innovation that consists of learnable activation functions, with the potential to capture more complex relationships from data. Presently, KANs are deployed by replacing multilayer perceptrons (MLPs) in deep networks, including advanced architectures such as vision Transformers (ViTs). Given the success of replacing MLP with KAN, this work asks whether KAN could learn token interactions. In this paper, we design the first learnable attention called Kolmogorov-Arnold Attention (KArAt) for ViTs that can operate on any basis, ranging from Fourier, Wavelets, Splines, to Rational Functions. However, learnable activations in the attention cause a memory explosion. To remedy this, we propose a modular version of KArAt that uses a low-rank approximation. By adopting the Fourier basis into this, Fourier-KArAt and its variants, in some cases, outperform their traditional softmax counterparts, or show comparable performance on CIFAR-10, CIFAR-100, and ImageNet-1K datasets. We also deploy Fourier KArAt to ConViT and Swin-Transformer, and use it in detection and segmentation with ViT-Det. We dissect the performance of these architectures on the classification task by analyzing their loss landscapes, weight distributions, optimizer paths, attention visualizations, and transferability to other datasets, and contrast them with vanilla ViTs. KArAt’s learnable activation yields a better attention score across all ViTs, indicating improved token-to-token interactions and contributing to enhanced inference. Still, its generalizability does not scale with larger ViTs. However, many factors, including the present computing interface, affect the relative performance of parameter- and memory-heavy KArAts. We note that the goal of this paper is not to produce efficient attention or challenge the traditional activations; by designing KArAt, we are the first to show that attention can be learned and encourage researchers to explore KArAt in conjunction with more advanced architectures that require a careful understanding of learnable activations. Our open-source code and implementation details are available on: subhajitmaity.me/KArAt
Keywords Multi-Head Self-Attention $\cdot$ Vision Transformers $\cdot$ Kolmogorov-Arnold Network $\cdot$ Kolmogorov-Arnold Transformers $\cdot$ Kolmogorov-Arnold Attention
@@ 1 Introduction
Figure 1: (a) Model parameters vs. Top-1 accuracy in ImegeNet-1K training of vanilla ViTs dosovitskiy2020vit , Vision KAN (DeiT+KAN) by VisionKAN2024 , ViT+KAN and Kolmogorov-Arnold Transformer (KAT) by yang2024kolmogorov .(b-i) The traditional softmax attention. (b-ii) The Kolmogorov-Arnold Attention (KArAt) replaces the softmax with a learnable operator, $\Phi^{i,j}$. (b-iii) Regular KArAt uses an operator matrix, $\Phi^{i,j}$ with $N^{2}$ learnable units acting on each row of ${\cal A}^{i,j}$, and is prohibitively expensive. (b-iv) Modular KArAt uses an operator $\widehat{\Phi}^{i,j}\in\mathbb{R}^{N\times r}$ with $r\ll N$, followed by a learnable linear projector $W\in\mathbb{R}^{r\times N}$.
@@
Artificial general intelligence has become a rapidly growing research direction, and Kolmogorov-Arnold Network (KAN) liu2024kan2 marks a remarkable innovation in that. KANs with learnable activation functions can potentially capture more complex relationships and facilitate meaningful interaction between the model and human intuition. After training a KAN on a specific problem, researchers can extract the learned univariate functions that the model uses to approximate complex multivariable functions. By studying these learned functions, researchers can gain insights into the underlying relationships from the data and refine the model. KANs exhibit state-of-the-art performance in finding symbolic function representations yu2024kanmlpfairercomparison , continual learning of one-dimensional functions liu2024kan .
@@
KANs were integrated with neural network architectures, the primary ones being conventional MLPs or convolution neural networks (CNNs). E.g., ferdaus2024kanice ; bodner2024convolutional ; drokin2024kolmogorov ; abueidda2025deepokan ; wang2025intrusion combine KAN with CNNs, li2024UKAN combine KAN with U-Net, etc. Interestingly, (yu2024kanmlpfairercomparison, ) claimed to make the first fairer comparison between KANs and MLPs on multiple ML tasks on small-scale datasets. By setting the same parameter counts in both networks, yu2024kanmlpfairercomparison concluded that KANs with B-Spline basis significantly underperform compared to their MLP counterparts in all ML tasks except symbolic formula representation; azam2024suitability reinstates a similar claim for vision. However, KAN’s exploration of more advanced architectures, such as Transformers, remains limited. VisionKAN2024 replace the MLP layers inside the encoder blocks of data-efficient image Transformers (DeiTs) deit with KANs and proposed DeiT+KAN, yang2024kolmogorov proposed two variants: ViT+KAN that replaces the MLP layers inside ViT’s encoder blocks, and Kolmogorov-Arnold Transformer (KAT), albeit similar to ViT+KAN but with a refined group-KAN strategy.
@@
While KANs were claimed to be not superior to simple MLPs in vision tasks yu2024kanmlpfairercomparison ; azam2024suitability , we were curious to find out how ViTs can see images through the lens of KANs. In that effort, Figure 1(a) demonstrates the performance of vanilla ViTs (-Tiny, -Small wu2022tinyvit ; steiner2022howtotrainvit , and -Base dosovitskiy2020vit ) and their KAN counterparts on ImageNet-1K deng2009imagenet training. While different variants of DeiT+KAN VisionKAN2024 and ViT+KAN yang2024kolmogorov show a 5%–18% drop in the Top-1 accuracy compared to vanilla ViTs and DeiTs in ImageNet-1K, the KATs yang2024kolmogorov show approximately 1%–2.5% gain in Top-1 accuracy compared to vanilla ViTs and DeiTs by keeping about the same number of parameters. These results show KANs require further investigation in more advanced architectures such as Transformers. Nevertheless, first, we want to understand why there is a discrepancy in these two seemingly equivalent implementations.
@@
Searching for an answer, we realized that while DeiT+KAN and ViT+KAN replace the MLP layer with KAN in the Transformer’s encoder block, KAT implements a sophisticated group-KAN strategy that reuses a learnable function among a group of units in the same layer and chooses different bases for different encoder blocks. This strategy improves accuracy by keeping almost the same parameter counts. Therefore, simply replacing MLPs with KANs might not guarantee better performance, but a properly designed KAN could. But, can we only learn attention?
Irrespective of the tasks or datasets at hand, traditional multihead attention, the heart of the Transformers, employs a softmax non-linearity as a one-size-fits-all mechanism for token interactions without the necessary justification for its use. We hypothesize that adopting learnable KAN operators could be a better fit for approximating the underlying non-linear relationship between image patches, providing the flexibility for modeling token interactions in ViTs. Therefore, in the rise of next-generation Transformers, such as Show-o xie2025showo , RL-VLM-F wang2024rlvlmf , Google’s TITAN behrouz2024titans , and SAKANA AI’s Transformer2 transformer2 , that mimic the human brain, we ask: How can we deploy learnable multi-head self-attention to the (vision) Transformers?
@@
We note that there is a line of research that proposes efficient attention mechanisms in terms of sparse attention fibottention ; yun2020n ; shi2021sparsebert ; kovaleva2019BERT ; zhang2021multi ; zaheer2020big ; guo2019star or linear/kernelized attention hanbridging ; nguyen2023probabilistic ; nguyen2021fmmformer ; nguyen2022fourierformer ; lu2021soft ; nguyen2022improving-icml , to remedy computational and memory complexities of attention calculation. In contrast, this paper tries to understand how learnable multi-head self-attention (MHSA) modules can perform over regular self-attention used in ViTs, a defining technology in vision in the last six years. Taken together, our contributions are:
@@
Designing a Learnable Attention Module for ViTs. This pilot study is on the self-attention in ViTs. In §3, we propose a general learnable Kolmogorov-Arnold multi-head self-attention, or KArAt, for ViTs that can operate on any basis. Due to the memory explosion, any full-rank learnable MHSA operator in ideal KArAt causes a training bottleneck. We propose a low-rank approximation of the operator to reduce memory requirements to a feasible level, but leave the possibility of exploring alternative techniques. By adopting the Fourier basis into this modular version, we design Fourier KArAt; see §4 for details. We cannot perform tasks with varying sequences and cross-attention without major modifications to KArAt. The most prominent modalities other than vision, such as audio, speech, video, and text, have varying sequence length, involve modeling with cross-attention-based encoder-decoder architectures, for which token interactions would be substantially different from self-attention. These are non-trivial technical challenges for implementing KArAt.
@@
Benchmarking, Evaluation, and Analysis (§56). We benchmark Fourier KArAt on CIFAR-10, CIFAR-100, krizhevsky2009cifar10/100 and ImageNet-1K deng2009imagenet datasets and evaluate their performance against vanilla ViTs (ViT-Tiny, -Small wu2022tinyvit ; steiner2022howtotrainvit , and -Base dosovitskiy2020vit ), and other popular ViTs (ConViT convit and Swin-Transformer liu2021swin ). We use small-scale datasets, SVHN svhn , Oxford Flowers 102 flowers102 , and STL-10 stl10 for understanding the transfer learning capability of the models. Additionally, we embed KArAt with ViT-Det vitdet for the object detection and segmentation on MS COCO coco . We dissect KArAts’ performance and generalization capacity by analyzing their loss landscapes, weight distributions, optimizer path, and attention visualization, and compare them with traditional softmax attention in vanilla ViTs. In addition, we investigate KArAts’ explanability for visual understanding; see §5.3.
@@ 2 Background
In MLP, the nonlinear activation functions, $\Phi(\cdot)$, are fixed bengio2017deep . For supervised tasks, given a training dataset $D$ with $N$ elements of the form (input, ground-truth) pairs, $\{(x_{i},y^{\star}_{i})\}_{i=1}^{N}$, the loss, $\mathcal{L}(X,Y^{\star}|\mathcal{W})=d_{Y}(a^{[L]},Y^{\star})$ (metric induced by the space $\mathbb{R}^{d_{L}}$) is calculated in the forward pass. The layerwise weights, $\{(W_{l},b_{l})\}_{l\in[L]}$ are learned by minimizing $\mathcal{L}(X,Y^{\star}|\mathcal{W})$.
Kolmogorov-Arnold Network (KAN) liu2024kan is a neural network involving learnable activations parametrized by a chosen set of basis functions defined on a set of grid points or knots. The idea for this network stems from the Kolmogorov-Arnold Representation Theorem.
@@
kolmogorov1956 For any multivariate continuous function, $f:[0,1]^{n}\to\mathbb{R}$, there exists a finite composition of continuous single-variable functions, $\phi_{q,p}:[0,1]\to\mathbb{R},\Phi_{q}:\mathbb{R}\to\mathbb{R}$ such that $f(x)=f(x_{1},x_{2},\cdots,x_{n})=\sum_{q=1}^{2n+1}\Phi_{q}\left(\sum_{p=1}^{n}\phi_{q,p}(x_{p})\right).$
Theorem 1 describes an exact finite expression using 2 layers, but liu2024kan generalized this representation to multi-layers via KAN; it is analogous to the Universal Approximation Theorem hornik1989multilayer . For input $x^{[0]}$, an $L$-layer KAN is given as $\text{KAN}(x^{[0]})=\Phi^{L}\circ\cdots\circ\Phi^{l}\circ\cdots\circ\Phi^{1}(x^{[0]})$. In each KAN layer, the activation function $\Phi$ is learnable, and $\Phi^{[l]}=[\Phi_{ij}^{[l]}]$} operates on $x^{[l-1]}\in\mathbb{R}^{d_{l-1}}$} to produce $x^{[l]}\in\mathbb{R}^{d_{l}}$}, such that:
@@
Originally, liu2024kan chose the B-spline basis functions. That is, the activation functions were defined to be $\phi(x)=w(b(x)+\text{spline}(x))$, where $b(x)=\texttt{SiLU}(x)=\frac{x}{1+e^{-x}}$ and $\text{spline}(x)=\sum\limits_{i}c_{i}B_{i}(x)$, with $B_{i}(\cdot)$ being one of the $k$-th degree B-spline basis functions parametrized by the $G$ grid points on a uniform grid, [I,I][-I,I]. In this case, the representation for each function is given as:
$).\Phi_{ij}^{[l]}(x^{[l-1]}_{j})=w^{[l]}_{ij}(\texttt{SiLU}(x^{[l-1]}_{j})+\sum\limits_{m=1}^{G+k-1}c^{[l]}_{ijm}B_{ijm}(x^{[l-1]}_{j}$)).
The weights $\{([c^{[l]}_{ijm}],[w^{[l]}_{ij}])\}_{l=1}^{L}$} are learned by minimizing the loss for a supervised learning task. The bases for KANs’ activations could be Fourier mehrabian2024implicit ; xu2024fourierkan , wavelet bozorgasl2024wav , fractals yang2024kolmogorov , etc.
Multi-head self-attention (MHSA) in ViTs dosovitskiy2020vit . The encoder-only vanilla ViT architecture is inspired by the Transformer proposed in attention . For simplicity, we do not mention the details of the layer normalization and other technicalities. Our central focus is the MHSA architecture. For limited space, we move this discussion to §A.
@@ 3 How Can We Design Learnable Attention?
Last year, we witnessed a surge in embedding KANs in different DNN architectures genet2024tkan ; ferdaus2024kanice ; li2024UKAN . To our knowledge, the work closely related to ours is KAT, where yang2024kolmogorov replaces the MLP layers in the ViTs with KAN layers; also, see VisionKAN2024 . We note that, for attentive graph neural networks (GNNs), fang2025kaa unifies the scoring functions of attentive GNNs and names it as Kolmogorov-Arnold Attention (KAA). However, it is orthogonal to our MHSA design; deploying learnable MHSA is complicated. In this Section, we design learnable multi-head self-attention (MHSA) module for ViTs, with improved interpretability, adaptability, and expressiveness that operate inside its encoder blocks.
@@
Projection onto the probability simplex. The softmax function acts on each row of $\cal A$} to create a probability distribution; the learnable attention does not guarantee that. To ensure that each row vector of $\Phi(\mathcal{A})$ lies on a probability simplex, we project them onto the $\ell_{1}$-unit ball. That is, for each attention matrix, ${\cal A}^{i,j}\in\mathbb{R}^{N\times N}$, and for each $k\in[N],$ we want to have $\|\tilde{\sigma}\left[({\cal A}^{i,j}_{k,:})\right]\|_{1}=1$} and $\tilde{\sigma}\left[({\cal A}^{i,j}_{k,l})\right]\geq 0$ for $l\in[N]$. We cast a sparse approximation problem, whose variants have been well-studied in the past decade and arise frequently in signal processing bryan2013making ; donoho2006compressed ; dutta_thesis and matrix approximation boas2017shrinkage ; stewart1993early ; see §3. Algorithm 1 provides an $\ell_{1}$-projection pseudocode; for each $k\in[N]$, setting $y=\tilde{\sigma}\left[({\cal A}^{i,j}_{k,:})\right]$} and $m=N$ in Algorithm 1, we obtain the projected attention vector.